OpenStudy (anonymous):

the drive wheel of a car is 38 inches in a diameter . how many times will it turn over a distance of 1km.?

OpenStudy (compphysgeek):

again you need the conversion factor from inches to meter. if you've got that you can calculate the circumference of the wheel in meters. Once you know that you can divide 1000 m by the circumference and you'll get the number of turns

OpenStudy (anonymous):

How many πD s are there in 1km 1 inch = 0.0254m 38 inch = 0.9652 m πD = 0.9652π 1km = 1000m N = 1000 / 0.9652π = 330 revolutions.

OpenStudy (anonymous):

correct?

OpenStudy (compphysgeek):

that's correct

OpenStudy (anonymous):

any solution that can i cancel out?

OpenStudy (compphysgeek):

you've got the first step \[ 38 in \times \frac{0.0254 m}{1 in} \] I dont see anything else where you can cancel anything out

OpenStudy (compphysgeek):

ok .. the last step of course \[ \frac{1000 m}{0.9652 m \times \pi} \] that gives you a number without any units

OpenStudy (anonymous):

329.7864...

OpenStudy (compphysgeek):

yes, rounded 330 as you had before

OpenStudy (anonymous):

what is km?

OpenStudy (compphysgeek):

what do you mean?

OpenStudy (anonymous):

330 is distance?

OpenStudy (compphysgeek):

no, 330 is the number of revolutions. It doesn't have a unit

OpenStudy (anonymous):

ah ok

OpenStudy (compphysgeek):

when we divided 1000m by \(0.96xx m \times \pi\) the meters cancelled out.